# Refusing Safe Prompts for Multi-modal Large Language Models

### v1

@@ Abstract
Multimodal large language models (MLLMs) have become the cornerstone of today’s generative AI ecosystem, sparking intense competition among tech giants and startups. In particular, an MLLM generates a text response given a prompt consisting of an image and a question. While state-of-the-art MLLMs use safety filters and alignment techniques to refuse unsafe prompts, in this work, we introduce MLLM-Refusal, the first method that induces refusals for safe prompts. In particular, our MLLM-Refusal optimizes a nearly-imperceptible refusal perturbation and adds it to an image, causing target MLLMs to likely refuse a safe prompt containing the perturbed image and a safe question. Specifically, we formulate MLLM-Refusal as a constrained optimization problem and propose an algorithm to solve it. Our method offers competitive advantages for MLLM model providers by potentially disrupting user experiences of competing MLLMs. Since competing MLLM’s users will receive unexpected refusals when they unwittingly use these perturbed images in their prompts. We evaluate MLLM-Refusal on four MLLMs across four datasets, demonstrating its effectiveness in causing competing MLLMs to refuse safe prompts while not affecting non-competing MLLMs. Furthermore, we explore three potential countermeasures—adding Gaussian noise, DiffPure, and adversarial training. Our results show that they are insufficient: though they can mitigate MLLM-Refusal’s effectiveness, they also sacrifice the accuracy and/or efficiency of the competing MLLM. The code is available at https://github.com/Sadcardation/MLLM-Refusal.
@@ 1 Introduction
Countermeasures. To counter the refusal perturbations created by MLLM-Refusal, we evaluate three countermeasures: Gaussian noise, DiffPure [39], and adversarial training. Gaussian noise reduces MLLM-Refusal’s effectiveness but significantly lowers MLLM accuracy, from 0.92 to around 0.80 with standard deviation $\sigma$=0.02. DiffPure uses a diffusion model to purify images, reducing MLLM-Refusal’s refusal rates but also dropping accuracy from 0.92 to 0.78 and increasing inference time by up to 13.07%. Adversarial training, which involves fine-tuning an MLLM on images with refusal perturbations, reduces refusal rates to around 60% but also significantly decreases MLLM accuracy and requires substantial computational resources. These findings indicate that while these countermeasures can mitigate MLLM-Refusal, they also sacrifice MLLM utility and efficiency.
@@ 2.2 Adversarial Examples
Adversarial examples are intentionally crafted or perturbed inputs causing a machine learning model to make incorrect predictions [51]. Adversarial perturbations are often nearly-imperceptible to humans but can significantly affect a model’s behavior. For MLLMs, adversarial examples can be applied to the image [47, 4, 43, 34, 7, 9, 20] and/or question [2, 24] in a prompt. In this work, we focus on perturbations that are added to an image in a normal user’s prompt to an MLLM. This is because a user may obtain the image from an untrusted source (e.g., Internet), in which an attacker may add a perturbation to it. Additionally, it is more challenging for an attacker to perturb the question in a normal user’s prompt since it is often came up by the user himself/herself.
@@ 3.1 System Setup
Our system setup involves three key roles: a model provider, competitors, and normal users. The model provider and competitors are resourceful companies, such as Meta, Google, and OpenAI, which invest substantially in training and deploying their own MLLMs as online chatbots or APIs. These companies fiercely compete for MLLM market dominance, continually enhancing their models’ capabilities and user experiences. We refer to a competitor’s MLLM as a competing MLLM and assume normal users query MLLMs with safe prompts.
@@ 4.1 Overview
Figure 3 is an overview of our MLLM-Refusal. Generally speaking, given an image, our MLLM-Refusal aims to optimize a refusal perturbation that achieves all effectiveness, locality, and stealthiness goals when added to the image. First, we construct a set of shadow questions using an LLM such as GPT-4, based on the model provider’s knowledge on exact, similar, or general user questions. Second, our MLLM-Refusal optimizes the refusal perturbation so that competing MLLMs are highly likely to refuse prompts with the perturbed image and shadow questions. We hypothesize that refusals to actual user questions stem from the transferability between shadow questions and actual user questions. Since different MLLMs have distinct vision-language projectors and we do not optimize the refusal perturbation for the model provider’s MLLM, the refusal perturbation is unlikely to impact the model provider’s MLLM. This means that the locality goal is naturally achieved. Finally, to ensure stealthiness of the refusal perturbation, our MLLM-Refusal applies $\ell_{\infty}$-norm constraints during the optimization.
@@ 5.2 Experimental Results
MLLM-Refusal outperforms compared methods:  Table 3 shows the refusal rates of compared methods and our MLLM-Refusal using three types of shadow questions, with LLaVA-1.5 as the competing MLLM on the VQAv2 dataset. We make four key observations. First, MLLM-Refusal consistently achieves the highest refusal rates across all shadow question types. Specifically, MLLM-Refusal achieves a refusal rate of 0.88, while Bagdasaryan et al. achieve a refusal rate of only 0.51 when the shadow questions are general user questions. Second, Qi et al. [43] is ineffective when extended to our scenarios, achieving near-zero refusal rates. This ineffectiveness arises because their method uses only images and does not incorporate shadow questions. Consequently, the refusal perturbations optimized with an empty set of shadow questions are unlikely to cause refusals for the competing MLLM when users query with actual non-empty questions. This indicates the necessity of constructing a set of shadow questions to optimize effective refusal perturbations. Third, the refusal rates are zero when no perturbations are added to images, demonstrating that the image-question pairs in the VQAv2 dataset are safe prompts to LLaVA-1.5, as they do not cause refusals.
Fourth, MLLM-Refusal + PGD can achieve comparable refusal rates to MLLM-Refusal. However, MLLM-Refusal + PGD requires more iterations to converge, resulting in higher computation costs, as shown in Table 4. Therefore, we use MLLM-Refusal in the following experiments.
@@ 6 Countermeasures
Images with refusal perturbation crafted by MLLM-Refusal can be considered as a type of adversarial example. Various countermeasures [39, 18, 8, 32] have been proposed to defend against adversarial examples. We categorize these countermeasures into testing-time and training-time countermeasures. Our MLLM-Refusal is evaluated against two popular testing-time countermeasures: Gaussian noise and DiffPure [39], as well as one training-time countermeasure: adversarial training [18].
@@
Figure 17: Accuracy of the competing MLLM and refusal rates of MLLM-Refusal when adding Gaussian noise. We use three types of shadow questions with LLaVA-1.5 on VQAv2. RR stands for refusal rate.
@@
Figure 17 shows the accuracy of the competing MLLM on visual question answering and the refusal rates of MLLM-Refusal when adding Gaussian noise as the countermeasure. The competing MLLM is LLaVA-1.5, and the dataset is VQAv2. We have two main observations. First, larger noise (i.e., larger $\sigma$) added to the image inputs is more effective at mitigating the impact of MLLM-Refusal. For example, without Gaussian noise (i.e., $\sigma=0$), the refusal rates of MLLM-Refusal with three types of shadow questions are all higher than 0.90. When $\sigma=0.02$, the refusal rates of MLLM-Refusal with three types of shadow questions are nearly zero. Second, adding Gaussian noise compromises the accuracy of the competing MLLM on visual question answering. Specifically, the accuracy drops significantly from 0.92 to around 0.80 when adding Gaussian noise with $\sigma=0.02$. Therefore, adding Gaussian noise as a countermeasure is insufficient.
Figure 18: Accuracy of the competing MLLM and refusal rates of MLLM-Refusal when using DiffPure. We use three types of shadow questions with LLaVA-1.5 on VQAv2.
@@
Figure 18 shows the accuracy of the competing MLLM and the refusal rates of MLLM-Refusal when using DiffPure with different timesteps. We utilized three types of shadow questions with LLaVA-1.5 on VQAv2. Similar to adding Gaussian noise, DiffPure reduces the effectiveness of MLLM-Refusal but significantly compromises the accuracy of the competing MLLM. Specifically, one timestep in DiffPure reduces the refusal rates of MLLM-Refusal from above 0.90 to below 0.20, while the accuracy decreases from 0.92 to 0.82. With two timesteps, the refusal rates drop further to near zero, and the accuracy decreases from 0.82 to 0.78.
@@ 7 Discussion and Limitations
Multi-round visual question answering:  As the capability to process increasingly lengthy contexts becomes critical for MLLMs, users can pose a series of successive questions regarding the initial image in multi-round question answer with an MLLM. In this case, we consider one multi-round question answering session where only initial prompt containing an image with refusal perturbation, the following question answering takes previous question-answer history in context. We want to study how many round an image with refusal perturbation can impact during multi-round question answering. Figure 20 shows the refusal rate in multi-round question answering for different types of shadow questions. For the reason of simplicity, the same user question is used in all chatting rounds. We observe that when the shadow questions use exact or similar user questions—the refusal rate tends to stabilize at a relatively high level as the number of chatting rounds increases, with a reduction in refusal rate of less than 5%. Conversely, longer context lengths reduce the effectiveness of added perturbations when shadow questions are general user questions, causing the refusal rate to decrease from 88% to around 65%. An interesting future work is to enhance the robustness of our MLLM-Refusal for multi-round visual question answering under different types of shadow questions. For instance, we could incorporate historical context into the optimization process of perturbation and expand our shadow questions set for better generalization to user questions.
@@
More modalities:  MLLMs are expanding to incorporate more modalities, such as audio [46, 40] and video [46, 40, 28, 57, 36], alongside text and image. As MLLMs become increasingly sophisticated, handling ever more complex input modalities, the potential vulnerabilities for refusing safe prompts across these new modalities also grow. An interesting future work is to extend our MLLM-Refusal to these additional modalities. For instance, we could explore adding nearly-imperceptible perturbations to audio waveforms or specific frames within videos, causing advanced MLLMs to refuse safe prompts just as effectively as with images in this work.
@@ 8 Conclusion and Future Work
In this work, we introduce MLLM-Refusal, the first method to induce refusals for safe prompts in MLLMs. Our method optimizes a nearly-imperceptible refusal perturbation that, when added to an image, causes competing MLLMs to refuse safe prompts while not affecting non-competing models. We demonstrate MLLM-Refusal’s effectiveness and locality across four MLLMs and datasets, highlighting its potential to gain competitive advantages for the model provider via disrupting user experiences of competing MLLMs. Our evaluation of countermeasures, including Gaussian noise, DiffPure, and adversarial training, reveals their insufficiency in mitigating MLLM-Refusal: they significantly sacrifice accuracy or efficiency of a competing MLLM in order to mitigate MLLM-Refusal’s effectiveness. Future work includes extending our method to multi-round visual question answering to further improve effectiveness and other modalities.

### v2

@@ Abstract
Multimodal large language models (MLLMs) have become the cornerstone of today’s generative AI ecosystem, sparking intense competition among tech giants and startups. In particular, an MLLM generates a text response given a prompt consisting of an image and a question. While state-of-the-art MLLMs use safety filters and alignment techniques to refuse unsafe prompts, in this work, we introduce MLLM-Refusal, the first method that induces refusals for safe prompts. In particular, our MLLM-Refusal optimizes a nearly-imperceptible refusal perturbation and adds it to an image, causing target MLLMs to likely refuse a safe prompt containing the perturbed image and a safe question. Specifically, we formulate MLLM-Refusal as a constrained optimization problem and propose an algorithm to solve it. Our method offers competitive advantages for MLLM model providers by potentially disrupting user experiences of competing MLLMs, since competing MLLM’s users will receive unexpected refusals when they unwittingly use these perturbed images in their prompts. We evaluate MLLM-Refusal on four MLLMs across four datasets, demonstrating its effectiveness in causing competing MLLMs to refuse safe prompts while not affecting non-competing MLLMs. Furthermore, we explore three potential countermeasures—adding Gaussian noise, DiffPure, and adversarial training. Our results show that though they can mitigate MLLM-Refusal’s effectiveness, they also sacrifice the accuracy and/or efficiency of the competing MLLM. The code is available at https://github.com/Sadcardation/MLLM-Refusal.
@@ 1 Introduction
Countermeasures. To counter the refusal perturbations created by MLLM-Refusal, we evaluate three countermeasures: adding Gaussian noise, DiffPure [40], and adversarial training. Gaussian noise reduces MLLM-Refusal’s effectiveness but significantly lowers MLLM accuracy, from 0.92 to around 0.80 with standard deviation $\sigma$=0.02. DiffPure uses a diffusion model to purify images, reducing MLLM-Refusal’s refusal rates but also dropping accuracy from 0.92 to 0.78 and increasing inference time by up to 13.07%. Adversarial training, which involves fine-tuning an MLLM on images with refusal perturbations, reduces refusal rates to around 60% but also significantly decreases MLLM accuracy and requires substantial computational resources. These findings indicate that while these countermeasures can mitigate MLLM-Refusal, they also sacrifice MLLM utility and efficiency.
@@ 2.2 Adversarial Examples
Adversarial examples are intentionally crafted or perturbed inputs causing a machine learning model to make incorrect predictions [52]. Adversarial perturbations are often nearly-imperceptible to humans but can significantly affect a model’s behavior. For MLLMs, adversarial examples can be applied to the image [48, 4, 44, 35, 7, 9, 21, 59] and/or question [2, 25] in a prompt. In this work, we focus on perturbations that are added to an image in a normal user’s prompt to an MLLM. This is because a user may obtain the image from an untrusted source (e.g., Internet), in which an attacker may add a perturbation to it. Additionally, it is more challenging for an attacker to perturb the question in a normal user’s prompt since it is often came up by the user himself/herself.
@@ 3.1 System Setup
Our system setup involves three key roles: a model provider, competitors, and normal users. The model provider and competitors are resourceful companies, such as Meta, Google, and OpenAI, which invest substantially in training and deploying their own MLLMs as online chatbots or APIs. These companies fiercely compete for MLLM market dominance, continually enhancing their models’ capabilities and user experiences. We refer to a competitor’s MLLM as a competing MLLM and assume normal users query MLLMs with safe prompts. The model provider aims to design refusal perturbations for images such that a competing MLLM will refuse to answer a safe prompt that includes an image with refusal perturbation and an accompanying question.
@@
Refusal perturbation:  A refusal perturbation is a nearly-imperceptible modification added to an image, causing competing MLLMs to respond with refusal to safe prompts containing the perturbed image. We formulate the crafting of refusal perturbations as a constrained optimization problem, as detailed in Section 4.3.
@@ 4.1 Overview
Figure 3 is an overview of our MLLM-Refusal. Generally speaking, given an image, our MLLM-Refusal aims to optimize a refusal perturbation that achieves all effectiveness, locality, and stealthiness goals when added to the image. First, we construct a set of shadow questions using an LLM such as GPT-4, and based on the model provider’s knowledge they can be exact, similar, or general user questions. Second, our MLLM-Refusal optimizes the refusal perturbation so that competing MLLMs are highly likely to refuse prompts with the perturbed image and shadow questions. We hypothesize that refusals to actual user questions stem from the transferability between shadow questions and actual user questions. Since different MLLMs have distinct vision-language projectors and we do not optimize the refusal perturbation for the model provider’s MLLM, the refusal perturbation is unlikely to impact the model provider’s MLLM. This means that the locality goal is naturally achieved. Finally, to ensure stealthiness of the refusal perturbation, our MLLM-Refusal applies $\ell_{\infty}$-norm constraints during the optimization.
@@ 5.2 Experimental Results
Table 5: Accuracy of compared methods and MLLM-Refusal using three types of shadow questions, with LLaVA-1.5 as the competing MLLM on the VQAv2 dataset.
No Perturbation 0.92 0.92 0.92
Qi et al. [44] 0.48 0.48 0.48
Bagdasaryan et al. [4] 0.03 0.04 0.03
MLLM-Refusal+PGD 0.03 0.03 0.04
MLLM-Refusal 0.03 0.04 0.03
MLLM-Refusal outperforms compared methods:  Table 3 shows the refusal rates of compared methods and our MLLM-Refusal using three types of shadow questions, with LLaVA-1.5 as the competing MLLM on the VQAv2 dataset. We make four key observations. First, MLLM-Refusal consistently achieves the highest refusal rates across all shadow question types. Specifically, MLLM-Refusal achieves a refusal rate of 0.88, while Bagdasaryan et al. [4] achieve a refusal rate of only 0.51 when the shadow questions are general user questions. Second, Qi et al. [44] is ineffective when extended to our scenarios, achieving near-zero refusal rates. This ineffectiveness arises because their method uses only images and does not incorporate shadow questions. Consequently, the refusal perturbations optimized with an empty set of shadow questions are unlikely to cause refusals for the competing MLLM when users query with actual non-empty questions. This indicates the necessity of constructing a set of shadow questions to optimize effective refusal perturbations. Third, the refusal rates are zero when no perturbations are added to images, demonstrating that the image-question pairs in the VQAv2 dataset are safe prompts to LLaVA-1.5, as they do not cause refusals. Fourth, MLLM-Refusal + PGD can achieve comparable refusal rates to MLLM-Refusal. However, MLLM-Refusal + PGD requires more iterations to converge, resulting in higher computation costs, as shown in Table 4. Therefore, we use MLLM-Refusal in the following experiments.
Table 5 shows the accuracy of LLaVA-1.5 on the VQAv2 dataset for the compared methods and MLLM-Refusal using three types of shadow questions. We observe that LLaVA-1.5 achieves high accuracy on the original, unperturbed images. However, the accuracy drops significantly when refusal perturbations are applied. Qi et al.[44] reduce accuracy by almost half, while Bagdasaryan et al.[4] and our MLLM-Refusal cause the accuracy of the competing MLLM to drop to nearly zero. This demonstrates that refusal perturbations can effectively alter the MLLM’s understanding of perturbed images, leading to inaccurate responses.
@@ 6 Countermeasures
Images with refusal perturbation crafted by MLLM-Refusal can be considered as a type of adversarial example. Various countermeasures [40, 19, 8, 33] have been proposed to defend against adversarial examples. We categorize these countermeasures into testing-time and training-time countermeasures. Our MLLM-Refusal is evaluated against two popular testing-time countermeasures: Gaussian noise and DiffPure [40], and one training-time countermeasure: adversarial training [19].
@@
Figure 17(a) shows the accuracy of the competing MLLM on visual question answering and the refusal rates of MLLM-Refusal when adding Gaussian noise as the countermeasure. The competing MLLM is LLaVA-1.5, and the dataset is VQAv2. We have two main observations. First, larger noise (i.e., larger $\sigma$) added to the image inputs is more effective at mitigating the impact of MLLM-Refusal. For example, without Gaussian noise (i.e., $\sigma=0$), the refusal rates of MLLM-Refusal with three types of shadow questions are all higher than 0.90. When $\sigma=0.02$, the refusal rates of MLLM-Refusal with three types of shadow questions are nearly zero. Second, adding Gaussian noise compromises the accuracy of the competing MLLM on visual question answering. Specifically, the accuracy drops significantly from 0.92 to around 0.80 when adding Gaussian noise with $\sigma=0.02$. Therefore, adding Gaussian noise as a countermeasure is insufficient.
Figure 17: Accuracy and refusal rates (RR) of MLLM-Refusal with (a) adding Gaussian noise and (b) using DiffPure. We use three types of shadow questions with LLaVA-1.5 on VQAv2.
@@
Figure 17(b) shows the accuracy of the competing MLLM and the refusal rates of MLLM-Refusal when using DiffPure with different timesteps. We utilized three types of shadow questions with LLaVA-1.5 on VQAv2. Similar to adding Gaussian noise, DiffPure reduces the effectiveness of MLLM-Refusal but significantly compromises the accuracy of the competing MLLM. Specifically, one timestep in DiffPure reduces the refusal rates of MLLM-Refusal from above 0.90 to below 0.20, while the accuracy decreases from 0.92 to 0.82. With two timesteps, the refusal rates drop further to near zero, and the accuracy decreases from 0.82 to 0.78.
@@ 7 Discussion and Limitations
Multi-round visual question answering:  The capability to process increasingly lengthy contexts has becomes a critical aspect when evaluating MLLMs. In this case, we consider one multi-round question answering session where only initial prompt containing an image with refusal perturbation, and the following question answering takes previous question-answer history in context to study how history affects the effect of MLLM-Refusal. Figure 19 shows the refusal rate in multi-round question answering for different types of shadow questions. For the reason of simplicity, the same user question is used in all chatting rounds. When the shadow questions use exact or similar user questions, the refusal rate tends to stabilize at a relatively high level with less than a 5% reduction as the number of chatting rounds increases. Conversely, longer context lengths reduce the effectiveness of added perturbations in the case of general user questions, causing the refusal rate to decrease from 88% to around 70%. An interesting future work is to enhance the effectiveness of our MLLM-Refusal for multi-round visual question answering under different types of shadow questions. For instance, we could incorporate multi-round visual question answering into the generation process of refusal perturbation for better effectiveness of MLLM-Refusal against repeated user questions.
@@
More modalities:  MLLMs are expanding to incorporate more modalities, such as audio [47, 41] and video [47, 41, 29, 58, 37], alongside text and image. As MLLMs become increasingly sophisticated, handling more complex input modalities, the potential vulnerabilities for refusing safe prompts across these new modalities also grow. An interesting future work is to extend our MLLM-Refusal to these additional modalities. For instance, we could explore adding nearly-imperceptible perturbations to audio waveforms or specific frames within videos, causing advanced MLLMs to refuse safe prompts just as effectively as with images in this work.
Potential countermeasures:  One countermeasure for MLLM users is only using images from trusted sources, which are less likely to add refusal perturbations. However, in practice, defining a trusted source is challenging. For instance, Meta is generally considered as a trusted source by its product users. However, as discussed in Section 3, it is possible for Meta, as an MLLM provider, to add refusal perturbations to gain competitive advantages. Another type of potential countermeasure involves provably robust defenses. For example, randomized smoothing [14] could be extended for this purpose. Randomized smoothing aggregates the model’s multiple outputs for a given input with randomly added Gaussian noise, providing a robustness guarantee for the aggregated output when the perturbation on an image input is bounded by a threshold. However, an MLLM’s response differs from a classifier’s predicted label, making it an interesting future work to explore how to aggregate MLLM responses and extend randomized smoothing to defend against our MLLM-Refusal.
@@ 8 Conclusion and Future Work
In this work, we introduce MLLM-Refusal, the first method to induce refusals for safe prompts in MLLMs. Our method optimizes a nearly-imperceptible refusal perturbation that, when added to an image, causes competing MLLMs to refuse safe prompts while not affecting non-competing MLLMs. We demonstrate MLLM-Refusal’s effectiveness and locality across four MLLMs and datasets, highlighting its potential to gain competitive advantages for the model provider via disrupting user experiences of competing MLLMs. Our evaluation of countermeasures, including Gaussian noise, DiffPure, and adversarial training, reveals their insufficiency in mitigating MLLM-Refusal: they significantly sacrifice accuracy or efficiency of a competing MLLM in order to mitigate MLLM-Refusal’s effectiveness. Future work includes extending our method to multi-round visual question answering to further improve effectiveness and other modalities.